# An EPTAS for multiprocessor scheduling with rejection under a machine cost constraint

### v1

@@ 4.1 Data normalization
We fix an instance $I$ of $Pm\mid rej,U,r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$ for discussion. The first step in the EPTAS is to call Approx1 and by Theorem 6, it produces a feasible solution such that the cost $Z(I)$ satisfies $\frac{1}{2}Z(I)\leq Z^{*}(I)\leq Z(I)$. We normalize the instance by dividing each $p_{j},e_{j}$ by the quantity $Z(I)$, and thus without loss of generality, we assume
@@
We set up a fixed constant $0<\Delta<\epsilon$ in Eq.(9) based on the values of $m,\epsilon$. Since $\frac{1}{\epsilon}$ is an integer, $\frac{1}{\Delta}$ is an integer as well.
$\Delta=\frac{\epsilon}{4m+12}.$ (9)
We partition the jobs into risky and safe using the constant $\Delta$.
@@
A job $J_{j}$ is risky if $e_{j}\geq\Delta$ and otherwise, $J_{j}$ is safe.
Note that if $e_{j}<\Delta$, then the rejection penalty is relatively small and thus it is safe to reject it.
@@
At most $\frac{1}{\Delta}$ risky jobs are rejected in the optimal solution.
@@
The proof is trivial since otherwise, the optimal cost is greater than $1$, which contradicts Eq.(8). ∎
@@ 4.2 Restricted instance
In this section, we construct another instance $I^{\prime}$ from the input instance $I$ but the number of distinct rejection penalty becomes a fixed constant. To achieve the goal, we set up another fixed constant $0<\delta<\Delta<\epsilon$ in Eq.(10). We note that $\frac{1}{\delta}$ is an integer since $\frac{1}{\Delta}$ and $\frac{1}{\epsilon}$ are both integers.
$\delta=\frac{\epsilon\Delta}{2}=\frac{\epsilon^{2}}{8m+24}.$ (10)
@@
(Restricted instance $I^{\prime}$) The restricted instance $I^{\prime}$ is constructed from the input instance $I$ by rounding the rejection penalty of each risky job up to the least multiple of $\delta$.
We let $J^{\prime}_{j}$ be the job in $I^{\prime}$ corresponding to the job $J_{j}$ in $I$, whose processing time and rejection penalty are denoted as $p^{\prime}_{j}$ and $e^{\prime}_{j}$, respectively. By the construction of $I^{\prime}$, we have $p^{\prime}_{j}=p_{j}$ for every job $J_{j}$. If $J_{j}$ is safe, then $e^{\prime}_{j}=e_{j}$. Otherwise, $e_{j}\geq\Delta$ and thus, by Eq.(10), $e^{\prime}_{j}\leq e_{j}+\delta\leq(1+\frac{\epsilon}{2})e_{j}$ We conclude that for every job $J_{j}$, we have $e^{\prime}_{j}\leq(1+\frac{\epsilon}{2})e_{j}$. Moreover, we know $e^{\prime}_{j}\geq\Delta$ if and only if $e_{j}\geq\Delta$. In other words, $J^{\prime}_{j}$ is risky if and only if $J_{j}$ is risky. We conclude all those facts in the following remark.
@@
For the instances $I$ and $I^{\prime}$, we have the following three facts:
@@
    $p^{\prime}_{j}=p_{j}$ for every job $J_{j}$;
@@
    $e^{\prime}_{j}\leq(1+\frac{\epsilon}{2})e_{j}$ for every job $J_{j}$;
  • 3.
    $J_{j}$ is risky if and only if $J^{\prime}_{j}$ is risky.
@@
    At most $\frac{1}{\Delta}$ risky jobs are rejected;
@@
    the cost is at most $(1+\frac{\epsilon}{2})Z^{*}(I)$.
@@
Consider the optimal solution $\pi^{*}$ for the instance $I$ and let $R^{*}$ denote the rejected job set. We construct $\pi$ from $\pi^{*}$ as follows: $J^{\prime}_{j}$ is rejected in $\pi$ if and only if $J_{j}$ is rejected in $\pi^{*}$. Once a job is accepted, then it will be scheduled on the same machine as in $\pi^{*}$. Clearly, $\pi$ is a feasible solution for $I^{\prime}$ since the processing times and the upper bound $U$ are not changed.
Note that we make the same rejection decision for each job as in $\pi^{*}$. By Lemma 8 and Statement 3 in Remark 10, at most $\frac{1}{\Delta}$ risky jobs are rejected and the first statement holds.
We next prove the second statement. By Statement 1 of Remark 10 and the construction of $\pi$, the makespan of $\pi$ is the same as $\pi^{*}$ and it is sufficient to show the total rejection penalty is bounded by $\sum_{J_{j}\in R^{*}}(1+\frac{\epsilon}{2})e_{j}$, which follows from Statement 2 in Remark 10. ∎
Note that in the restricted instance $I^{\prime}$, the rejection penalty of each risky job is a multiple of $\delta$ and is bounded by $1$. Therefore, we can partition the risky jobs into at most $\frac{1}{\delta}$ groups in the following definition.
@@
$B_{i}=\{J_{j}\in B:e_{j}=i\delta\},\mbox{ for every $i=1,2,\ldots,1/\delta$.}$
Since the jobs in $B_{i}$ have identical rejection penalties, we sort the jobs in $B_{i}$ by the non-increasing order of their processing times. This way, we have $J=S\cup\bigcup_{i=1}^{1/\delta}B_{i}$.
By Lemma 11 and Definition 12, at most $\frac{1}{\Delta}$ jobs in $\bigcup_{i=1}^{1/\delta}B_{i}$ are rejected.
@@
(Rejection types) A rejection type is a $\frac{1}{\delta}$-tuple $(k_{1},k_{2},\ldots,k_{1/\delta})$ such that each $k_{i}$ is a non-negative integer and $\sum_{i=1}^{1/\delta}k_{i}\leq 1/\Delta$.
A feasible solution of $I^{\prime}$ is associated with a rejection type $(k_{1},k_{2},\ldots,k_{1/\delta})$ if in the solution, exactly $k_{i}$ jobs of $B_{i}$ are rejected, for every $i=1,\ldots,1/\delta$.
@@
The number of distinct rejection types is at most $N_{r}=(1/\Delta+1)^{1/\delta}$, which is a fixed constant when $m$ is a constant.
@@
Since $\sum_{i=1}^{1/\delta}k_{i}\leq 1/\Delta$ and $k_{i}$ is a non-negative integer, we have $k_{i}\in\{0,1,\ldots,1/\Delta\}$ for every $i=1,\ldots,1/\delta$. The lemma is proved because $1/\delta$ and $1/\Delta$ are fixed integers. ∎
Note that the feasible solution $\pi$ stated in Lemma 11 must be associated with a unique rejection type.
@@
Suppose that $\pi$ is associated with the rejection type $(k_{1},k_{2},\ldots,k_{1/\delta})$. There exists a feasible solution $\pi^{\prime}$ for the restricted instance $I^{\prime}$ such that
@@
    $\pi^{\prime}$ is associated with $(k_{1},k_{2},\ldots,k_{1/\delta})$, the same as $\pi$;
@@
    the cost is at most $(1+\frac{\epsilon}{2})Z^{*}(I)$;
@@
Assume otherwise, in the solution $\pi$, there exists a pair of jobs $(p_{j},p_{t})$ in some job set $B_{i}$ such that $j\leq k_{i}\leq t$ and; $J_{j}$ is accepted and $J_{t}$ is rejected in $\pi$. By Definition 12, we have $e_{j}=e_{t}$ and $p_{j}\geq p_{t}$. We construct another solution $\pi_{1}$ from $\pi$ as follows: We accept $J_{t}$ and schedule it on the machine, where $J_{j}$ is assigned in $\pi$. We reject $J_{j}$ and for every job except $J_{j},J_{t}$, we make the same rejection decision and schedule it on the same machine as in $\pi$ if it is accepted. Clearly, the first item is satisfied.
Note that $p_{j}\geq p_{t}$ and thus $\pi_{1}$ is a feasible solution and by the construction of $\pi_{1}$, the makespan of $\pi_{1}$ is not greater than $\pi$. Since $e_{j}=e_{t}$, the rejection penalties of $\pi$ and $\pi_{1}$ are exactly the same. Therefore, the cost of $\pi_{1}$ does not exceed the cost of $\pi$. By Lemma 11, the second item in the lemma is satisfied. We can continue to ”swap” the rejection decisions of two jobs until the third item is satisfied, which leads to a desired solution. ∎
By Lemma 14, there are at most $N_{r}=(1/\Delta+1)^{1/\delta}$ distinct rejection types. Therefore, we are able to enumerate all rejection types to find the one associated with $\pi^{\prime}$ in Lemma 15. Given the rejection type $(k_{1},k_{2},\ldots,k_{1/\delta})$ of $\pi^{\prime}$, it is easy to determine the rejection decisions for all risky jobs. In more details, we reject the first $k_{i}$ jobs of $B_{i}$, for every $i=1,\ldots,1/\delta$ and accept the other risky jobs. We use the notations $A^{\prime}_{1},R^{\prime}_{1}$ to denote the accepted and rejected risky jobs, respectively.
Definition 16.
A job $J^{\prime}_{j}\in A^{\prime}_{1}$ is large if $p^{\prime}_{j}\geq\Delta$; otherwise, $J^{\prime}_{j}$ is tiny. Let $L$ and $T$ be the large and tiny job set and thus $A^{\prime}_{1}=L\cup T$.

### v2

@@ 4.1 Data normalization
We fix an instance $I$ of $P\mid rej,U,r_{i}=r\mid C_{\max}(A)+\sum_{J_{j}\in R}e_{j}$ for discussion. The first step in the PTAS is to call Approx1 and by Theorem 1, it produces a feasible solution such that the cost $Z(I)$ satisfies $\frac{1}{2}Z(I)\leq Z^{*}(I)\leq Z(I)$. We normalize the instance by dividing each $p_{j},e_{j}$ and $U$ by the quantity $Z(I)$, and thus without loss of generality, we assume
@@
We set up a fixed constant $0<\gamma<\epsilon$ in Eq.(9) based on the values of $\epsilon$. Since $\frac{1}{\epsilon}$ is an integer, $\frac{1}{\gamma}$ is an integer as well.
$\gamma=\frac{\epsilon}{24}.$ (9)
We partition the jobs into risky and safe (large and tiny, respectively) using the constant $\gamma^{3}$ ($\gamma$, respectively).
@@
A job $J_{j}$ is risky if $e_{j}\geq\frac{1}{2}\gamma^{3}$ and otherwise, $J_{j}$ is safe. Moreover, $J_{j}$ is large if $p_{j}\geq\gamma$ and otherwise, $J_{j}$ is tiny.
Note that if $e_{j}<\frac{1}{2}\gamma^{3}$, then the rejection penalty is relatively small and thus it is safe to reject it. One sees that if $e_{j}>1$, then by Eq.(8), $J_{j}$ must be accepted and if $p_{j}>1$, then it must be rejected in the optimal schedule. Therefore, we can immediately make the rejection decisions for those jobs and for simplicity, we assume that all $e_{j}$ and $p_{j}$ are at most $1$.
@@
At most $\frac{2}{\gamma^{3}}$ risky jobs are rejected and at most $\frac{1}{\gamma}$ large jobs are scheduled on a same machine in the optimal solution.
@@
If there are $\frac{2}{\gamma^{3}}+1$ risky jobs are rejected, then the optimal cost is greater than 11, which contradicts Eq.(8). Moreover, if at least $\frac{1}{\gamma}+1$ large jobs are scheduled on a same machine, then the makespan is greater than 11, a contradiction to Eq.(8) again. ∎
@@ 4.2 Restricted instance
In this section, we construct another instance $I^{\prime}$ from the input instance $I$ but the number of distinct rejection penalty becomes a fixed constant. To achieve the goal, we set up another fixed constant $0<\delta<\frac{1}{2}\gamma^{3}<\epsilon$ in Eq.(10). We note that $\frac{1}{\delta}$ is an integer since $\frac{1}{\epsilon}$ is an integer.
$\delta=\frac{\epsilon\gamma^{3}}{12}=\frac{\epsilon^{4}}{12\cdot 24^{3}}.$ (10)
@@
(Restricted instance $I^{\prime}$) The restricted instance $I^{\prime}$ is constructed from the input instance $I$ by rounding the rejection penalty of each risky job up to the least multiple of $\delta$  [^3] By Lemma 3, we can determine the rejection decisions of risky jobs in $O(n^{2/\gamma^{3}})$ time. But constructing the restricted instance helps us reduce the time to $O(1)$ (See $k_{i}$ in Definition 7 and the third item in Lemma 6).
One clearly sees that the processing time of each job is unchanged. For ease of presentation, we still denote by $\{J_{1},J_{2},\ldots,J_{n}\}$ the job set in $I^{\prime}$, where the processing time of $J_{j}$ is $p_{j}$ but its rejection penalty is denoted as $e^{\prime}_{j}$. If $J_{j}$ is safe in $I$, then $e^{\prime}_{j}=e_{j}$. Otherwise, $e_{j}\geq\frac{1}{2}\gamma^{3}$ and thus, by Eq.(10), $e^{\prime}_{j}\leq e_{j}+\delta\leq(1+\frac{\epsilon}{6})e_{j}$ We conclude that for every job $J_{j}$, we have $e^{\prime}_{j}\leq(1+\frac{\epsilon}{6})e_{j}$. Moreover, since $\frac{1}{2}\gamma^{3}$ is a multiple of $\delta$ by Eq.(10), we know $e^{\prime}_{j}\geq\frac{1}{2}\gamma^{3}$ if and only if $e_{j}\geq\frac{1}{2}\gamma^{3}$. In other words, $J_{j}$ is risky in $I^{\prime}$ if and only if $J_{j}$ is risky in $I$. We conclude all those facts in the following remark.
@@
For the instances $I$ and $I^{\prime}$, we have the following two facts:
@@
    $e^{\prime}_{j}\leq(1+\frac{\epsilon}{6})e_{j}$ for every job $J_{j}$;
@@
    $J_{j}$ is risky in $I^{\prime}$ if and only if $J_{j}$ is risky in $I$.
@@
    At most $\frac{2}{\gamma^{3}}$ risky jobs are rejected;
@@
    at most $\frac{1}{\gamma}$ large jobs are scheduled on a same machine;
  • 3.
    the cost is at most $(1+\frac{\epsilon}{6})Z^{*}(I)$.
@@
Consider the optimal solution $\pi^{*}$ for the instance $I$ and let $R^{*}$ denote the rejected job set. We construct $\pi$ from $\pi^{*}$ as follows: $J_{j}$ is rejected in $\pi$ if and only if $J_{j}$ is rejected in $\pi^{*}$. Once a job is accepted, then it will be scheduled on the same machine as in $\pi^{*}$. Clearly, $\pi$ is a feasible solution for $I^{\prime}$ since the processing times and the upper bound $U$ are not changed.
Note that we make the same rejection decision for each job as in $\pi^{*}$. By Lemma 3 and Statement 2 in Remark 1, the first and second statements hold.
We next prove the third statement. By the construction of $\pi$, the makespan of $\pi$ is the same as $\pi^{*}$ and it is sufficient to show the total rejection penalty is bounded by $\sum_{J_{j}\in R^{*}}(1+\frac{\epsilon}{6})e_{j}$, which follows from Statement 1 in Remark 1. ∎
Recall that all $e_{j}$ and $p_{j}$ are at most $1$ and $1$ is a multiple of $\delta$. Therefore, for each risky job $J_{j}$ in $I^{\prime}$, its rejection penalty $e^{\prime}_{j}$ is a multiple of $\delta$ and is bounded by $1$. So, we can partition the risky jobs into at most $\frac{1}{\delta}$ groups. For the jobs that are both safe and large, we partition them into $\frac{1}{\gamma^{2}}$ in the following definition.
@@
$B_{i}=\{J_{j}\in B:e^{\prime}_{j}=i\delta\},\mbox{ for every $i=1,2,\ldots,1/\delta$.}$
Since the jobs in $B_{i}$ have identical rejection penalties, we sort the jobs in $B_{i}$ by the non-increasing order of their processing times.
Moreover, we let $L$ and $T$ denote the set of large and tiny jobs in $I^{\prime}$ (See Definition 4). Note that the processing time of a large job is at most 11 by Eq.(8). Therefore, we partition the jobs in $S\cap L$ into the following subsets:
$SL_{i}=\{J_{j}\in S\cap L:i\gamma^{2}\leq p_{j}<(i+1)\gamma^{2}\},\mbox{ for every $i=1,\ldots,1/\gamma^{2}$.}$
Some sets of $B_{i},SL_{i}$ may be empty and this will not influence our analysis.
By Lemma 4 and Definition 6, at most $\frac{2}{\gamma^{3}}$ jobs in $\bigcup_{i=1}^{1/\delta}B_{i}$ are rejected and at most $\frac{m}{\gamma}$ jobs in $S\cap L=\bigcup_{i=1}^{1/\gamma^{2}}SL_{i}$ are accepted.
@@
(Rejection types) A rejection type is a $(\frac{1}{\gamma^{2}}+\frac{1}{\delta}+1)$-tuple $(\ell_{1},\ldots,\ell_{1/\gamma^{2}},k_{1},\ldots,k_{1/\delta},K)$ such that
  • 1.
    each $\ell_{i}$ is a non-negative integer and $\sum_{i=1}^{1/\gamma^{2}}\ell_{i}\leq m/\gamma$;
  • 2.
    each $k_{i}$ is a non-negative integer and $\sum_{i=1}^{1/\delta}k_{i}\leq 2/\gamma^{3}$;
  • 3.
    $K$ is a non-negative integer that is at most $\frac{2}{\gamma^{3}}+\frac{8}{\gamma^{2}}$.
A feasible solution of $I^{\prime}$ is associated with a rejection type $(\ell_{1},\ldots,\ell_{1/\gamma^{2}},k_{1},\ldots,k_{1/\delta},K)$ if in the solution, exactly $\ell_{i}$ jobs of $SL_{i}$ are accepted, for every $i=1,\ldots,1/\gamma^{2}$ and; exactly $k_{i}$ jobs of $B_{i}$ are rejected, for every $i=1,\ldots,1/\delta$ and; the total penalty of the rejected jobs in $S\cap T$ is in the interval $\frac{K}{2}\gamma^{3},\frac{K+1}{2}\gamma^{3})$, where $S\cap T$ is the set of jobs that are both safe and tiny.
@@
The number of distinct rejection types is at most $N_{r}=(\frac{m}{\gamma}+1)^{1/\gamma^{2}}\times(\frac{2}{\gamma^{3}}+1)^{1/\delta}\times(\frac{2}{\gamma^{3}}+\frac{8}{\gamma^{2}}+1)$, which is polynomial in $m$.
@@
Since $\sum_{i=1}^{1/\delta}\ell_{i}\leq m/\gamma$ and $\ell_{i}$ is a non-negative integer, we have $\ell_{i}\in\{0,1,\ldots,m/\gamma^{2}\}$ for every $i=1,\ldots,1/\gamma^{2}$. Therefore, we have at most $(\frac{m}{\gamma}+1)^{1/\gamma^{2}}$ distinct combinations of $\ell_{1},\ldots,\ell_{1/\gamma^{2}}$. The lemma is not hard to prove by discussing $k_{1},\ldots,k_{1/\delta}$ and $K$ separately. ∎
Note that the feasible solution $\pi$ stated in Lemma 4 has a cost at most $(1+\frac{\epsilon}{6})Z^{*}(I)$. Therefore, by Eq.(8), the rejected jobs in $S\cap T$ has a total penalty at most $1+\frac{\epsilon}{6}=(\frac{2}{\gamma^{3}}+\frac{8}{\gamma^{2}})\times\frac{1}{2}\gamma^{3}$. Therefore, $\pi$ must be associated with a rejection type.
@@
Suppose that $\pi$ is associated with the rejection type $(\ell_{1},\ldots,\ell_{1/\gamma^{2}},k_{1},\ldots,k_{1/\delta},K)$. There exists a feasible solution $\pi^{\prime}$ for the restricted instance $I^{\prime}$ such that
@@
    $\pi^{\prime}$ is associated with $(\ell_{1},\ldots,\ell_{1/\gamma^{2}},k_{1},\ldots,k_{1/\delta},K)$, the same as $\pi$;
@@
    the cost is at most $(1+\frac{\epsilon}{6})Z^{*}(I)$ and at most $\frac{1}{\gamma}$ large jobs are assigned to a same machine;
@@
Assume otherwise, in the solution $\pi$, there exists a pair of jobs $(p_{j},p_{t})$ in some job set $B_{i}$ such that $j\leq k_{i}\leq t$ and; $J_{j}$ is accepted and $J_{t}$ is rejected in $\pi$. By Definition 6, we have $e^{\prime}_{j}=e^{\prime}_{t}$ and $p_{j}\geq p_{t}$. We construct another solution $\pi_{1}$ from $\pi$ as follows: We accept $J_{t}$ and schedule it on the machine, where $J_{j}$ is assigned in $\pi$. We reject $J_{j}$ and for every job except $J_{j},J_{t}$, we make the same rejection decision and schedule it on the same machine as in $\pi$ if it is accepted. Clearly, the first item is satisfied.
Note that $p_{j}\geq p_{t}$ and thus $\pi_{1}$ is a feasible solution and by the construction of $\pi_{1}$, the makespan of $\pi_{1}$ is not greater than $\pi$. Since $e^{\prime}_{j}=e^{\prime}_{t}$, the rejection penalties of $\pi$ and $\pi_{1}$ are exactly the same. Therefore, the cost of $\pi_{1}$ does not exceed the cost of $\pi$. By Lemma 4, the second item in the lemma is satisfied. We can continue to ”swap” the rejection decisions of two jobs until the third item is satisfied, which leads to a desired solution. ∎
By Lemma 5, there are at most $N_{r}=(\frac{m}{\gamma}+1)^{1/\gamma^{2}}\times(\frac{1}{\gamma^{3}}+1)^{1/\delta}\times(\frac{2}{\gamma^{3}}+\frac{8}{\gamma^{2}}+1)$ distinct rejection types. Therefore, we are able to enumerate all rejection types to find the one associated with $\pi^{\prime}$ in Lemma 6. Given the rejection type $(\ell_{1},\ldots,\ell_{1/\gamma^{2}},k_{1},\ldots,k_{1/\delta},K)$ of $\pi^{\prime}$, it is easy to determine the rejection decisions for all risky jobs, i.e., the jobs in $B$. In more details, we reject the first $k_{i}$ jobs of $B_{i}$, for every $i=1,\ldots,1/\delta$ and accept the other risky jobs.
4.3 Determine the rejection decisions for safe jobs
In this section, suppose that we are given the rejection type $(\ell_{1},\ldots,\ell_{1/\gamma^{2}},k_{1},\ldots,k_{1/\delta},K)$ associated with $\pi^{\prime}$ stated in Lemma 6 and we will present a linear program to determine the rejection decision for every safe job. For every safe job $J_{j}\in S$, let $x_{j}\in\{0,1\}$ be a binary decision variable such that $x_{j}=1$ if and only if $J_{j}$ is rejected in the solution $\pi^{\prime}$. Moreover, we let $y_{j}=1-x_{j}$. We are now ready to present our linear program and similarly to LP(2), we relax $x_{j},y_{j}$ to be continuous.
$\min$ $\sum_{J_{j}\in S}y_{j}p_{j}$ (11)
$x_{j}+y_{j}=1$ $\forall J_{j}\in S$
$\sum_{J_{j}\in SL_{i}}y_{j}=\ell_{i}$ $\forall i=1,\ldots,1/\gamma^{2}$
$\sum_{J_{j}\in S\cap T}x_{j}e^{\prime}_{j}<\frac{K+1}{2}\gamma^{3}$
$\sum_{J_{j}\in S\cap T}x_{j}e^{\prime}_{j}\geq\frac{K}{2}\gamma^{3}.$
We next interpret the LP(11) in details. The first constraint is due to the definition of $y_{j}$ and the second indicates the number of the accepted jobs in $SL_{i}$ is exactly $\ell_{i}$. The third and fourth constraints are from the definition of $K$. We minimize the total processing time of the accepted safe jobs. Note that LP(11) has exactly $2|S|\leq 2n$ variables and $|S|+\frac{1}{\gamma^{2}}+2\leq n+\frac{1}{\gamma^{2}}+2$ constraints. The linear program is polynomial-time solvable in $n,m$ because the time complexity of a linear program is polynomial in the number of continuous variables and polynomial in the number of constraints [15]. Using the interior point method, the LP(11) can be solved in $O(n^{3.5}L)$ where $L$ is the size of the input.
We consider a basic solution of LP(11) and partition the safe jobs into three sets as follows.
$S_{1}=\{J_{j}\in S:\mbox{ $y_{j}=1$}\},S_{2}=\{J_{j}\in S:\mbox{ $x_{j}=1$}\},\mbox{ and }S_{3}=S\setminus(S_{1}\cup S_{2}).$ (12)